In one roll of two number cubes, what is the probability that the sum of the faces is 11?
Found 2 solutions by funmath, checkley71:
Answer by funmath(2933)

There are 2 ways to get an 11: (5,6) and (6,5)
There are 6*6=36 outcomes when rolling two dice.
Therefore the probability of getting an 11 is: 2/36
This can be reduced to:1/18
Happy Calculating!!!

Answer by checkley71(8403)

THE NUMBER OF WAYS TO GET AN 11 IS A 5&6 OR 6&5 THUS THERE ARE 2 OUT OF 36 POSSIBLE COMBINATIONS OR 2/36=1/18 OR 5.5555555%